OpenStudy (anonymous):

If I buy 2 apples and 6 oranges, I will spend $5.60. However, if I buy 6 apples and 2 oranges, I will spend $4.80. How much does one orange cost? A. $0.55 B. $0.70 C. $0.75 D. $0.60

OpenStudy (anonymous):

Lte A for apples, G for orangs And write equation 2A +6G =$5.6, (tim*(-3)) 6A +2G =$4.8 you can use elemate to solve or Subtitution I like solve by elemate -6A-18G=-$16.8 6A +2G = $4.8 ..................Add -16G = -$12 G = -$12/-16 G =$0.75 Subtitution G =$0.75 solve or A 2A +6(0.75) =$5. 2A +4.5=$5.6 2A=5.6-4.5 A = 1.1/2 A = $0.55 apples:$0.55 for each orang: $0.75 for each